OpenStudy (anonymous):

solve for X please. -12x+9=(-9x)+15

OpenStudy (anonymous):

idk stop asking questions so that they can help me and not you

OpenStudy (anonymous):

hahha theres alot of ppl they can help you too kyeteee mijaa

OpenStudy (anonymous):

-12x+9x=15-9 3x=6 x=6/3=2 thereforex=2

OpenStudy (anonymous):

@leah14 ^^^^^^ shhhhhhhhhh its not my fault they replyin to mee @Parvathysubhash thanks!!

OpenStudy (anonymous):

nope that was wrong @Parvathysubhash

OpenStudy (anonymous):

wc atleast giv me a medal

OpenStudy (anonymous):

noooooooooo

OpenStudy (anonymous):

ya i put it as my answer and got it wrongg

OpenStudy (anonymous):

no yaar its corrrecttt

OpenStudy (anonymous):

First add 12x to the other side. 9 = 3x + 15 you should be able to solve it from there!

OpenStudy (anonymous):

Subtract 15 from both sides. 6 = 3x Then divide both by 3. So x should equal 2. @Parvathysubhash got it right!

OpenStudy (anonymous):

wierd it said i got it wrong...

OpenStudy (anonymous):

ysss srry

OpenStudy (anonymous):

lols;

OpenStudy (anonymous):

somebody gave me a medall yaaay